Both Bond A and Bond B have 6.2 percent coupons and are priced at par value. Bond A has 6 years to maturity, while Bond B has 15 years to maturity.

a. If interest rates suddenly rise by 1 percent, what is the percentage change in price of Bond A and Bond B? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

%Δ in Price
Bond A %
Bond B (8.99) %

b. If interest rates suddenly fall by 1 percent instead, what would be the percentage change in price of Bond A and Bond B? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places.)

%Δ in Price
Bond A %
Bond B 10.25 %
Answer #1

1
=PV(6.2%+1%,6,-6.2%*1000,-1000)/1000-1=-4.73725176364753%

2
=PV(6.2%+1%,15,-6.2%*1000,-1000)/1000-1=-8.9939674811324%

3
=PV(6.2%-1%,6,-6.2%*1000,-1000)/1000-1=5.04338869943015%

4
=PV(6.2%-1%,15,-6.2%*1000,-1000)/1000-1=10.2407512760635%